# The Role of Video Generation in Enhancing Data-Limited Action Understanding

### v1

@@ 4.2.1 Zero-shot Action Recognition
Table 1 shows the zero-shot action recognition results of the proposed method. The proposed method has significant improvements across all three datasets, especially the HMDB-51, where the Top-1 accuracy of zero-shot action recognition is improved by 5% after combining the generated sample training. It should be noted that when our method is based solely on the VIT-B backbone, our performance on HMDB-51 even surpasses those of methods based on the VIT-L backbone ( zhu2023orthogonal ; akbari2023alternating ). The rich contextual information related to actions is introduced by the proposed method, leading to the improvement of zero-shot action recognition.
@@ 4.2.4 Long-Tail Action Recognition
The concept of “long-tail problem” refers to the phenomenon where the imbalance in the distribution of samples across different classes leads to a significantly larger number of samples in the minority head classes compared to the tail classes, as models tend to perform well on the head classes but exhibit poorer performance on the tail classes. Following perrett2023use , we construct long-tail action recognition datasets based on UCF-101 and HMDB-51.
@@ 4.2.5 Abnormal Action Detection
Video anomaly action detection aims to identify abnormal events or actions within videos. The primary objective is to pinpoint the specific time window when an anomalous activity occurs, such as crimes, traffic accidents, or other illegal activities, etc. This task is challenging because weak labels are provided only at the video level, whereas the model needs to make frame-level predictions for abnormal actions. We employ the MIL (Multiple Instance Learning) baseline method following lv2023unbiased for abnormal action detection. Initially, we first pre-train the model on generated samples and subsequently fine-tune it on real samples. We report the AUC and AUCA lv2021localizing of the frame-level ROC results on the UCF-Crime dataset.
@@ 4.3 Ablation Studies
The results show that incorporating either environmental information or character information in the synthetic samples improves zero-shot accuracy. The information enhancement strategy, which combines environmental information and character information, can lead to a 3.4% improvement compared to the basic strategy. This indicates that enhancing the action-related information of the generated samples is beneficial for training action understanding models. Figure LABEL:fig:IEsamples illustrates the visualization of the “brush hair” action samples generated using the four aforementioned strategies. When employing the basic strategy, the video generation model tends to generate similar videos in which the characters consistently appear as Caucasian women with long brown hair. When the “Env” strategy is adopted, various contextual details that assist in understanding the ”brush hair” action are included in generated videos, such as full-length mirrors and bathrooms with tiled backgrounds. This information enhances the knowledge encompassed by the generated dataset. We set the information enhancement strategy as the default setting for the remaining experiments.
@@ 5 Analysis
We conducted qualitative analysis on synthetic datasets and real datasets, as shown in Figure LABEL:fig:clust. We extract CLIP features from the HMDB-51 dataset, the synthetic HMDB-51 dataset with the basic strategy, and the synthetic HMDB-51 dataset with the information enhancement strategy. Subsequently, we make the t-SNE visualization of these datasets based on the extracted features.
In Figure LABEL:fig:clust, it is evident that the boundaries between different categories of synthetic samples are quite distinct. In contrast, the boundaries between samples of different categories in the real dataset are more blurred and lack clarity. This could be attributed to the higher complexity of real samples compared to synthetic samples. This complexity stems from various aspects such as changes in lighting, camera angles, and background details, among others, leading real samples to contain more information. When generating samples based on the basic strategy, the text2video model tends to produce videos with similar content, resulting in low levels of effective information contained in the generated datasets. This limits the upper bound of knowledge that the synthetic data can contain.
The proposed information enhancement strategy effectively increases the information content in the generated samples. As illustrated in Figure LABEL:fig:clust, the proposed method improves the information contained in the generated samples, bringing their distribution patterns closer to real samples.
We perform quantitative analyzes in Table 8. Davies-Bouldin (DB) is a metric used to assess the effectiveness of clustering algorithms, where a smaller value indicates clearer separation between clusters and tighter intra-cluster cohesion. We employ the Davies-Bouldin score to measure the complexity of the datasets. The first column of Table 8 presents the Davies-Bouldin scores for the generated datasets and real datasets. Real datasets (UCF-101,HMDB-51) exhibit higher DB scores compared to generated samples due to their complexity. The Maximum Mean Discrepancy(MMD) gretton2006kernel is a commonly used metric for measuring the domain gap between different datasets, where a higher MMD indicates a larger domain gap. In columns 2, 3, and 4 of Table 8, we present the MMD between various datasets and the HMDB-51 dataset. MMD between real datasets, such as UCF-101 and HMDB-51, exhibit lower values, while the MMD between generated datasets and real datasets are higher, indicating a larger domain gap. The proposed strategy reduces the domain gap between generated samples and real samples, which is manifested as lower MMD values.

### v2

@@ 4.2.1 Zero-shot Action Recognition
Table 1 shows the zero-shot action recognition results of the proposed method. The proposed method has significant improvements across all three datasets, especially the HMDB-51, where the Top-1 accuracy of zero-shot action recognition is improved by 5% after combining the generated sample training. It should be noted that when our method is based solely on the VIT-B backbone, our performance on HMDB-51 even surpasses those of methods based on the VIT-L backbone ( Zhu et al. (2023); Akbari et al. (2023)). The rich contextual information related to actions is introduced by the proposed method, leading to the improvement of zero-shot action recognition.
@@ 4.2.4 Long-Tail Action Recognition
The concept of “long-tail problem” refers to the phenomenon where the imbalance in the distribution of samples across different classes leads to a significantly larger number of samples in the minority head classes compared to the tail classes, as models tend to perform well on the head classes but exhibit poorer performance on the tail classes. Following Perrett et al. (2023), we construct long-tail action recognition datasets based on UCF-101 and HMDB-51.
@@ 4.2.5 Abnormal Action Detection
Video anomaly action detection aims to identify abnormal events or actions within videos. The primary objective is to pinpoint the specific time window when an anomalous activity occurs, such as crimes, traffic accidents, or other illegal activities, etc. This task is challenging because weak labels are provided only at the video level, whereas the model needs to make frame-level predictions for abnormal actions. We employ the MIL (Multiple Instance Learning) baseline method following Lv et al. (2023) for abnormal action detection. Initially, we first pre-train the model on generated samples and subsequently fine-tune it on real samples. We report the AUC and AUCA Lv et al. (2021) of the frame-level ROC results on the UCF-Crime dataset.
@@ 4.3 Ablation Studies
Figure 3: Visualization of generated samples. The strategies adopted from left to right are: “Basic”, “Env”, “Cha” and “IE”.
@@
The results show that incorporating either environmental information or character information in the synthetic samples improves zero-shot accuracy. The information enhancement strategy, which combines environmental information and character information, can lead to a 3.4% improvement compared to the basic strategy. This indicates that enhancing the action-related information of the generated samples is beneficial for training action understanding models. Figure 3 illustrates the visualization of the “brush hair” action samples generated using the four aforementioned strategies. When employing the basic strategy, the video generation model tends to generate similar videos in which the characters consistently appear as Caucasian women with long brown hair. When the “Env” strategy is adopted, various contextual details that assist in understanding the ”brush hair” action are included in generated videos, such as full-length mirrors and bathrooms with tiled backgrounds. This information enhances the knowledge encompassed by the generated dataset. We set the information enhancement strategy as the default setting for the remaining experiments.
@@ 5 Analysis
We conducted qualitative analysis on synthetic datasets and real datasets, as shown in Figure 1. We extract CLIP features from the HMDB-51 dataset, the synthetic HMDB-51 dataset with the basic strategy, and the synthetic HMDB-51 dataset with the information enhancement strategy. Subsequently, we make the t-SNE visualization of these datasets based on the extracted features.
In Figure 1, it is evident that the boundaries between different categories of synthetic samples are quite distinct. In contrast, the boundaries between samples of different categories in the real dataset are more blurred and lack clarity. This could be attributed to the higher complexity of real samples compared to synthetic samples. This complexity stems from various aspects such as changes in lighting, camera angles, and background details, among others, leading real samples to contain more information. When generating samples based on the basic strategy, the text2video model tends to produce videos with similar content, resulting in low levels of effective information contained in the generated datasets. This limits the upper bound of knowledge that the synthetic data can contain.
The proposed information enhancement strategy effectively increases the information content in the generated samples. As illustrated in Figure 1, the proposed method improves the information contained in the generated samples, bringing their distribution patterns closer to real samples.
We perform quantitative analyzes in Table 8. Davies-Bouldin (DB) is a metric used to assess the effectiveness of clustering algorithms, where a smaller value indicates clearer separation between clusters and tighter intra-cluster cohesion. We employ the Davies-Bouldin score to measure the complexity of the datasets. The first column of Table 8 presents the Davies-Bouldin scores for the generated datasets and real datasets. Real datasets (UCF-101,HMDB-51) exhibit higher DB scores compared to generated samples due to their complexity. The Maximum Mean Discrepancy(MMD) Gretton et al. (2006) is a commonly used metric for measuring the domain gap between different datasets, where a higher MMD indicates a larger domain gap. In columns 2, 3, and 4 of Table 8, we present the MMD between various datasets and the HMDB-51 dataset. MMD between real datasets, such as UCF-101 and HMDB-51, exhibit lower values, while the MMD between generated datasets and real datasets are higher, indicating a larger domain gap. The proposed strategy reduces the domain gap between generated samples and real samples, which is manifested as lower MMD values.